OpenStudy (anonymous):

How do i write a recursive formula for the explicit formula? A(n)=5+(n-1)(-7). Answer Choices: 1.A(n)=A(n-1)+5;A(1)=-7 2.A(n) A(n-1)-7;A(1)=5 3.A(n-1)=A(n)-7;A(1)=5 4.A(n-1)=A(n)+5; A(1)=-7 Which one should i choose?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

I need to help with figuring out the correct answer. what is the first steps i take to solve this problem?

OpenStudy (loser66):

You typed the wrong equation, I guess.

OpenStudy (anonymous):

Can you explain why that's the correct answer?

OpenStudy (loser66):

It should be \(A_n = 5+A_{n-1}(-7)\)

OpenStudy (anonymous):

No, on my homework its says what i typed on here correctly.

OpenStudy (loser66):

Can you scan and post the original one?

OpenStudy (loser66):

for recursive and explicit formula, need know the right one to solve.

OpenStudy (anonymous):

I don't have a scanner

OpenStudy (anonymous):

I can draw the formula.

OpenStudy (loser66):

ok, the very first formula, not the answer choices.

OpenStudy (anonymous):



OpenStudy (anonymous):

This is exactly what it looks like

OpenStudy (anonymous):

I'm sorry that's the best i can draw it or show to you. It is a explicit formula so it has paratheses not powers.

OpenStudy (loser66):

\(A_n = 5 +(n-1)(-7)\) if n=, then \(A_1= 5 +(1-1)(-7) =5\) ok?

OpenStudy (loser66):

if n=1 sorry about that :)

OpenStudy (loser66):

Need you confirm whether you understand step 1 or not

OpenStudy (anonymous):

I understand.

OpenStudy (loser66):

so, we have only 2 options b or c, right?

OpenStudy (anonymous):

no i have 4 answer options

OpenStudy (anonymous):

It could be B your right.

OpenStudy (loser66):

but for 1and 4, it gives us \(A_1=-7\) hence get rid of them, right?

OpenStudy (anonymous):

so i slash out all the other answers?

OpenStudy (loser66):

sure

OpenStudy (anonymous):

Okay that makes sense.

OpenStudy (loser66):

now, between 2, 3 we test one by one. If you are lucky enough, you get the answer at the first test

OpenStudy (anonymous):

okay how do you do that?

OpenStudy (anonymous):

Explain.

OpenStudy (loser66):

for 2) it says \(A_n = A_{n-1} -7\) and \(A_1=5\) right?

OpenStudy (anonymous):

Yes.

OpenStudy (loser66):

now, if n =2, then n-1 =1, so \(A_2= A_1 -7= 5 -7 =-2\) ok?

OpenStudy (anonymous):

Okay.

OpenStudy (loser66):

Let test \(A_n = 5 + (n-1) (-7) \\ A_2= 5 +(2-1)(-7) = -2\) Ah ha!!

OpenStudy (anonymous):

Is that the correct answer?

OpenStudy (loser66):

one more!! if n=3? apply the second option: \(A_n = A_{n-1}-7\\A_3=A_2-7 = -2-7=-9\) Apply the given equation: \(A_n = 5+(n-1)(-7)\\A_3= 5+(3-1)(-7)=5+2(-7) =5-14=-9\) Wow... it works, right? That is the way you work to figure out what it the right answer

OpenStudy (anonymous):

where did you get n=3?

OpenStudy (loser66):

just apply n=1, 2, 3, ..... to get arbitrary \(n^{th}\) term.

OpenStudy (loser66):

this is the order of the sequence, like you have a line of people and you name the first one is A1, the second one is A2, the third one is A3.....

OpenStudy (loser66):